OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1). A. y + 1 = -2(x – 5) B. y – 1 = -2(x – 5) C. y – 5 = -2(x + 1) D. y -5 = -2(x – 1)

OpenStudy (anonymous):

You need to do the same you did in the other question. Can you do it?

OpenStudy (anonymous):

\[y-y_0=m(x-x_0)\]

OpenStudy (anonymous):

y=-2x+9

OpenStudy (anonymous):

You know that m = -2

OpenStudy (anonymous):

so its y + 1 = -2(x – 5)

OpenStudy (anonymous):

Yeah, Dashni gave you the answr.

OpenStudy (anonymous):

Dhashni is right.